### v1

@@ 3 Robust computationally efficient Gauss-Newton policy optimization
Using MPC as policy within RL can significantly alleviate some of the main challenges of RL based on NN policies: it provides a good initial policy and can deal with hard constraints in real-time by solving an optimization problem. Although the sparse parameterization of MPC policies enables the use of second-order optimization methods, the training still remains computationally challenging, as Hessian estimates such as (20) rely on the second-order NLP sensitivities. For that reason, we propose in Section 3.1 a novel Gauss-Newton approximation of the deterministic policy Hessian that retains superlinear convergence while avoiding the computation of the second-order NLP sensitivities. We then present in Section 3.2 a new approach to introduce momentum in the Gauss-Newton update to increase numerical stability.
@@ 3.1 Gauss-Newton approximation of the deterministic policy Hessian
To be a unique maximum, $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ must be strictly negative definite. Since $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ can also be locally constant for $\boldsymbol{a}$ in the neighborhood of $\pi^{\star}(\boldsymbol{s})$ if the optimal policy is not unique, the curvature can also be zero. As the result $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ is negative semi-definite. $\qed$
Kordabad et al. (2022) proves that the approximation of the deterministic policy Hessian stated in (20) converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$. We show that even when neglecting the first part $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ of the approximation in (20), $\boldsymbol{M}_{2}(\boldsymbol{\theta})\approx\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})$ converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$ with the benefit that the second-order NLP sensitivities can be omitted. A Hessian approximation with the structure of $\boldsymbol{M}_{2}(\theta)$ is also referred to as a Gauss-Newton approximation (Bottou et al., 2018). We first state the following assumption.
@@
as it follows from Theorem 1 that $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ vanishes at $\boldsymbol{\theta}^{\star}$ and $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is assumed to be bounded.
@@
Here, $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}):\mathcal{S}\rightarrow[0,\infty)$ denotes the probability density that state $\boldsymbol{s}$ is encountered under closed-loop operation of $\pi_{\boldsymbol{\theta}}$. Since $\rho_{\pi_{\boldsymbol{\theta}}}$ and $\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$ can be uniformly bounded, the dominated convergence theorem (Bronštejn et al., 2016) can be applied so the limit operator can be moved inside the integral. It follows that
@@
where the equality in the last line holds because of (3.1), proving that $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ vanishes if $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})$ is bounded.
@@
Since $\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ is negative semi-definite for $\boldsymbol{\theta}$ in the neighborhood of $\boldsymbol{\theta}^{\star}$, it follows from the fact that the result of a left and right multiplication with any matrix and its transposed to a negative semi-definite matrix is again a negative semi-definite matrix (Bronštejn et al., 2016) so
@@
Since a unique optimal policy is assumed, $\boldsymbol{M}_{2}(\theta^{\star})=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ is strictly negative definite. Further $\boldsymbol{B}_{k}=\boldsymbol{M}_{2,k}$ converges to $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ as shown in Theorem 1. Therefore condition (38) is satisfied. $\qed$
3.2 Robust Gauss-Newton policy optimization with momentum
If the action-value function $Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ of the current policy $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is explicitly known, and the expected values could be computed perfectly, the proposed approximate Newton method would converge superlinearly. However, in reality, the action-value function of the current policy must be approximated. Further, the expected value must be computed empirically by drawing a finite amount of samples. Both can lead to noisy estimates of the policy gradient $\tilde{\boldsymbol{g}}_{k}\approx\boldsymbol{g}_{k}$ and the approximate policy Hessian $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$ preventing accurate parameter updates and causing potential instability. A common choice to mitigate the effect of noisy estimates within first-order stochastic optimization is the use of momentum-based approaches as introduced in Section 2.2.1. While second-order updates are usually robust with respect to poorly estimated Hessian approximations (Bollapragada et al., 2019; Bottou et al., 2018), the estimates still need to be of reasonable accuracy. A single poor Hessian estimation can already cause instability, for example if the eigenvalues that are estimated are orders of magnitude wrong, causing an update that deteriorates the parameter iterates to values that are far from optimal. We show next how the ideas of momentum can be transferred to second-order approaches by applying some adaptations.
Analogous to the update rule of stochastic gradient ascent with momentum (17), we introduce the update rule
$\boldsymbol{p}_{k}=-\alpha\hat{\boldsymbol{D}}_{k}^{-1}\hat{\boldsymbol{m}}_{k}$ (39)
in alignment with the subsampled Newton update rule (23). Here, $\hat{\boldsymbol{m}}_{k}$ denotes the unbiased averaged gradient estimate as described by the momentum-based updates in (15) and (16). The unbiased averaged Hessian estimate that we derive in this work is denoted by $\hat{\boldsymbol{D}}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$. Lets define the biased averaged Hessian estimate $\boldsymbol{D}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ recursively as
@@ 3.2 Robust Gauss-Newton policy optimization with momentum
with $0\leq\eta<1$ denoting the exponential averaging factor for Hessian estimation, $\boldsymbol{D}_{k-1}$ denoting the biased averaged Hessian estimation of the previous iteration, and $\tilde{\boldsymbol{B}}_{k}$ being the local Hessian estimate at the current iteration. Note that we do not specify $\tilde{\boldsymbol{B}}_{k}$ as it can be any Hessian approximation such as $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{k}$ or $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{2,k}$.
As in the case of the gradient momentum, the matrix $\boldsymbol{D}_{0}$ must be initialized. Although it may be appealing to initialize with a zero matrix $\boldsymbol{D}_{0}=0$, it turns out to be impractical because especially in the first few steps of the optimization, $\boldsymbol{D}_{k}$ would be severely influenced by inaccurate local estimates $\tilde{\boldsymbol{B}}_{k}$. Consider for instance the scenario that the eigenvalues $\lambda(\tilde{\boldsymbol{B}}_{1})$ of $\tilde{\boldsymbol{B}}_{1}$ are in the order of magnitude of $\lambda_{i}(\tilde{\boldsymbol{B}}_{1})\sim 1$ whereas the eigenvalues of $\boldsymbol{B}_{1}$ are rather in the order of magnitude of $\lambda_{i}(\boldsymbol{B}_{1})\sim 10^{3}$, the inverse of the Hessian in (39) would scale the gradient roughly by the inverse of the eigenvalues, which is $1/\lambda_{i}(\tilde{\boldsymbol{B}}_{1})\sim 1$ although it should rather be $1/\lambda_{i}(\boldsymbol{B}_{1})\sim 10^{-3}$. To avoid such undesirable behavior, we propose to initialize with a multiple of the identity matrix
$\boldsymbol{D}_{0}=-\omega^{-1}\boldsymbol{I}$ (41)
where $\omega>0$ can be understood as an initial learning rate since $\boldsymbol{D}_{0}^{-1}=-\omega\boldsymbol{I}$ would scale each element of the gradient by $\omega$ when applied in (39). Further, $\omega^{-1}$ should be chosen as a rough estimate of the largest absolute eigenvalue of $\boldsymbol{B}_{1}$.
To account for the initialization bias towards $\boldsymbol{D}_{0}$, we derive a scaling function similar to (16) that allows fast adjustments to newly obtained local Hessian estimates $\tilde{\boldsymbol{B}}_{k}$ and still maintains the added robustness from $\boldsymbol{D}_{0}$. Writing the recursive definition (40) explicitly and using $-\omega^{-1}\boldsymbol{I}=(1-\eta)\tilde{\boldsymbol{B}}_{0}$ gives
$\boldsymbol{D}_{k}=$ $-\eta^{k}\omega^{-1}\boldsymbol{I}+(1-\eta)\sum_{i=1}^{k}\eta^{k-i}\tilde{\boldsymbol{B}_{i}}$
@@
This expression aligns closely to results from Kingma and Ba (2015) with the exception that the sum starts with $i=0$ instead of $i=1$. Following a similar derivation as in Kingma and Ba (2015), we derive the following corrected exponential moving average $\hat{\boldsymbol{D}}_{k}$
@@
We show empirically in Section 5 that the proposed bias correction improves the Hessian estimate in the early stages and that the limit (44) is also observed in practice.
@@ 4 Analytical case study
$\nabla_{a}^{2}Q^{\pi_{\theta}}(s,a)=$ $2\gamma p_{\theta}-1$ (52b)
@@ 5 Continuously stirred tank reactor
We demonstrate the improved convergence behavior and data-efficiency of the proposed second-order Gauss-Newton approach in comparison with the state-of-the-art first-order optimizer Adam (Kingma and Ba, 2015), and show that the convergence behavior is similar to more elaborate but more costly approximate Hessian methods like (20) at the example of a continuously stirred tank reactor (CSTR). We also investigate the effect of the introduced momentum for the Hessian and derive tuning rules that we validate with empirical investigations. All code is available online [^1] https://github.com/DeanBrandner/GaussNewtonRLforMPC.

### v2

@@ 3 Robust computationally efficient Gauss-Newton policy optimization
Using MPC as policy within RL can significantly alleviate some of the main challenges of RL based on NN policies: it provides a good initial policy and can deal with hard constraints in real-time by solving an optimization problem. Although the sparse parameterization of MPC policies enables the use of second-order optimization methods, the training still remains computationally challenging, as Hessian estimates such as (22) rely on the second-order NLP sensitivities. For that reason, we build on the findings in kordabadQuasiNewtonIterationDeterministic2022 by deriving an approximation for the Hessian matrix that avoids the computation of the second-order NLP sensitivities. We derive and propose a novel Gauss-Newton approximation of the deterministic policy Hessian that retains superlinear convergence while avoiding the computation of the second-order NLP sensitivities in Section 3.1. We then present a new approach to introduce momentum in the Gauss-Newton Hessian estimate to increase numerical stability in Section 3.2 before we combine both findings in the policy optimization algorithm with adaptive trust-region in Section 3.3.
@@ 3.1 Gauss-Newton approximation of the deterministic policy Hessian
To be a unique maximum, $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ must be strictly negative definite. Since $Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})$ can also be locally constant for $\boldsymbol{a}$ in the neighborhood of $\pi^{\star}(\boldsymbol{s})$ if the optimal policy is not unique, the curvature can also be zero. Hence, $\nabla_{\boldsymbol{a}}^{2}Q^{\pi^{\star}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi^{\star}(\boldsymbol{s})}$ is negative semi-definite. $\qed$
kordabadQuasiNewtonIterationDeterministic2022 proves that the approximation of the deterministic policy Hessian stated in (22) converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$. The derived Hessian approximation obtains a superlinear convergence guarantee. However if MPC is used as a policy, it still requires to compute the second-order NLP sensitivities due to (23a), which can be computationally demanding. Building up on the findings in kordabadQuasiNewtonIterationDeterministic2022, we show that even when neglecting the first part $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ of the approximation in (22), $\boldsymbol{M}_{2}(\boldsymbol{\theta})\approx\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})$ converges to the true deterministic policy Hessian for $\boldsymbol{\theta}\rightarrow\boldsymbol{\theta}^{\star}$ with the benefit that the second-order NLP sensitivities can be omitted. A Hessian approximation with the structure of $\boldsymbol{M}_{2}(\theta)$ is also referred to as a Gauss-Newton approximation (bottouOptimizationMethodsLargeScale2018). We first state the following assumption.
@@
as it follows from Lemma 1 that $\nabla_{\boldsymbol{a}}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ vanishes at $\boldsymbol{\theta}^{\star}$ and $\nabla_{\boldsymbol{\theta}}^{2}\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is assumed to be bounded.
@@
Here, $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s}):\mathcal{S}\rightarrow[0,\infty)$ denotes the unnormalized probability density that state $\boldsymbol{s}$ is encountered under closed-loop operation of $\pi_{\boldsymbol{\theta}}$. Since $\rho_{\pi_{\boldsymbol{\theta}}}$ and $\varphi_{\boldsymbol{\theta}}(\boldsymbol{s})$ can be uniformly bounded, the dominated convergence theorem (bronstejnTaschenbuchMathematik2016) can be applied so the limit operator can be moved inside the integral. It follows that
@@
where the equality in the last line holds because of (3.1), proving that $\boldsymbol{M}_{1}(\boldsymbol{\theta})$ vanishes if $\rho_{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s})$ is bounded. $\qed$
@@
Since $\nabla_{\boldsymbol{a}}^{2}Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})|_{\boldsymbol{a}=\pi_{\boldsymbol{\theta}}(\boldsymbol{s})}$ is negative semi-definite for $\boldsymbol{\theta}$ in the neighborhood of $\boldsymbol{\theta}^{\star}$, it follows from the fact that the result of a left and right multiplication with any matrix and its transposed to a negative semi-definite matrix is again a negative semi-definite matrix (bronstejnTaschenbuchMathematik2016) so
@@
Since a unique optimal policy is assumed, $\boldsymbol{M}_{2}(\theta^{\star})=\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ is strictly negative definite. Further $\boldsymbol{B}_{k}=\boldsymbol{M}_{2,k}$ converges to $\nabla_{\boldsymbol{\theta}}^{2}J(\boldsymbol{\theta})|_{\boldsymbol{\theta}=\boldsymbol{\theta}^{\star}}$ as shown in Theorem 1. Therefore condition (41) is satisfied. $\qed$
3.2 Robust Hessian estimation via momentum
If the action-value function $Q^{\pi_{\boldsymbol{\theta}}}(\boldsymbol{s},\boldsymbol{a})$ of the current policy $\pi_{\boldsymbol{\theta}}(\boldsymbol{s})$ is explicitly known, and the expected values could be computed perfectly, the proposed approximate Newton method would converge superlinearly. However, in reality, the action-value function of the current policy must be approximated. Further, the expected value must be computed empirically by drawing a finite amount of samples. Both can lead to noisy estimates of the policy gradient $\tilde{\boldsymbol{g}}_{k}\approx\boldsymbol{g}_{k}$ and the approximate policy Hessian $\tilde{\boldsymbol{B}}_{k}\approx\boldsymbol{B}_{k}$ preventing accurate parameter updates and causing potential instability. A common choice to mitigate the effect of noisy estimates within first-order stochastic optimization is the use of momentum-based approaches as introduced in Section 2.2.1. While second-order updates are usually robust with respect to poorly estimated Hessian approximations (bollapragadaExactInexactSubsampled2019; bottouOptimizationMethodsLargeScale2018), the estimates still need to be of reasonable accuracy. A single poor Hessian estimation can already cause instability, for example if the eigenvalues that are estimated are orders of magnitude wrong, causing an update that deteriorates the parameter iterates to values that are far from optimal. We show next how the ideas of momentum can be transferred to second-order approaches by applying some adaptations.
Analogously to the exponentially averaged gradient estimate $\boldsymbol{m}_{k}$, we define the biased exponentially averaged Hessian estimate $\boldsymbol{D}_{k}\in\mathbb{R}^{n_{\boldsymbol{\theta}}\times n_{\boldsymbol{\theta}}}$ recursively as
@@ 3.2 Robust Hessian estimation via momentum
where $0\leq\eta<1$ denotes the exponential averaging factor for Hessian estimation, $\boldsymbol{D}_{k-1}$ denotes the biased averaged Hessian estimation of the previous iteration, and $\tilde{\boldsymbol{B}}_{k}$ denotes the local Hessian estimate at the current iteration. Note that we do not specify $\tilde{\boldsymbol{B}}_{k}$ as it can be any Hessian approximation such as $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{k}$ or $\tilde{\boldsymbol{B}}_{k}=\tilde{\boldsymbol{M}}_{2,k}$.
As in the case of the gradient momentum, the matrix $\boldsymbol{D}_{0}$ must be initialized. Although it may be appealing to initialize with a zero matrix $\boldsymbol{D}_{0}=0$, it turns out to be impractical because especially in the first few steps of the optimization, $\boldsymbol{D}_{k}$ would be severely influenced by inaccurate local estimates $\tilde{\boldsymbol{B}}_{k}$. To avoid such undesirable behavior, we propose to initialize with the square root of the first estimate of the second momentum of the Adam optimizer, which is
$\boldsymbol{D}_{0}=-\mathrm{diag}\left(\sqrt{\hat{\boldsymbol{v}}_{1}}+\epsilon\right).$ (43)
To account for the initialization bias towards $\boldsymbol{D}_{0}$, we derive a scaling function similar to (16) that allows fast adjustments to newly obtained local Hessian estimates $\tilde{\boldsymbol{B}}_{k}$ and still maintains the added robustness from $\boldsymbol{D}_{0}$. Writing the recursive definition (42) explicitly and using $\mathrm{diag}\left(\sqrt{\hat{\boldsymbol{v}}_{1}}+\epsilon\right)=(1-\eta)\tilde{\boldsymbol{B}}_{0}$ gives
$\boldsymbol{D}_{k}=$ $-\eta^{k}\mathrm{diag}\left(\sqrt{\hat{\boldsymbol{v}}_{1}}+\epsilon\right)+(1-\eta)\sum_{i=1}^{k}\eta^{k-i}\tilde{\boldsymbol{B}_{i}}$
@@
This expression aligns closely to results from kingmaAdamMethodStochastic2015 with the exception that the sum starts with $i=0$ instead of $i=1$. Following a similar derivation as in kingmaAdamMethodStochastic2015, we derive the following corrected exponential moving average $\hat{\boldsymbol{D}}_{k}$
@@
3.3 Robust adaptive trust-region constrained policy optimization
Although second order approaches can lead to improved convergence behavior, they can also induce instabilities during training. A frequently encountered scenario is that the Hessian of the objective function is singular, which is the case if the objective function behaves locally linear. Even if regularization is applied to the Hessian, the computed (subsampled) Newton steps (25) can be very large.
Trust-region constrained optimization approaches account for such degenerate behavior. They constrain the proposed update to be within a trust region radius $\delta$ around the local iterate. This trust region radius governs how conservative the local update is: a small trust region ensures stable but slow improvement, whereas a large trust region encourages large but potentially unstable steps. Algorithms with an adaptive trust region $\delta=\delta_{k}$ aim at finding a compromise between updates with significant improvements but with a fair amount of stability. A common approach is to compare the predicted improvement of the local quadratic approximation with the truly observed new objective value at the new iterate.
In RL, such comparative approaches are impractical due to the stochasticity of optimization problem (9). In fact, the expected value of the cumulative reward (5) needs to be estimated empirically. As a consequence, the empirical estimate is noisy and would be compared to a deterministic model prediction. If the empirical estimate is insufficiently accurate due to its noise, the trust region radius $\delta_{k}$ would decrease towards zero as the RL algorithm cannot find a feasible improvement although the model prediction may be good. Therefore, we propose to update the trust region radius $\delta_{k}$ based on the 2-norm of the Adam update step (19) and embed it into the following trust-region constrained optimization problem
$\boldsymbol{p}_{k}=\arg\min_{\boldsymbol{p}}\quad$ $\boldsymbol{p}^{\top}\hat{\boldsymbol{m}}_{k}+\frac{1}{2}\boldsymbol{p}^{\top}\hat{\boldsymbol{D}}_{k}\boldsymbol{p}$ (47a)
$\mathrm{s.t.}\quad$ $\left\|\boldsymbol{p}\right\|_{2}\leq\alpha\left\|\frac{\hat{\boldsymbol{m}}_{k}}{\sqrt{\hat{\boldsymbol{v}}_{k}}+\epsilon}\right\|_{2}=\delta_{k}.$ (47b)
The proposed trust region constrained optimization problem with the adaptive trust region now allows to obtain large update steps if the gradients are large in magnitude and far away from the optimal solution, whereas the trust region is scaled down if the algorithm approaches a local optimum.
@@ 4 Analytical case study
$\nabla_{a}^{2}Q^{\pi_{\theta}}(s,a)=$ $2\gamma p_{\theta}-1.$ (55b)
@@ 5 Continuously stirred tank reactor
We demonstrate the scalability with respect to the number of parameters of the proposed adaptive trust region constrained Gauss-Newton approach in comparison to more elaborate but more costly approximate Hessian methods like (22) at the example of a continuously stirred tank reactor (CSTR). We also investigate the resilience of the proposed adaptive trust region constrained Gauss-Newton approach with respect to parameter scales. All investigations are compared to the Adam optimizer (kingmaAdamMethodStochastic2015). Lastly, the proposed approach is compared to deep RL policies trained with the TD3 algorithm (fujimotoAddressingFunctionApproximation2018) with respect to learning efficiency and final closed-loop performance. All code is available online [^1] https://github.com/DeanBrandner/GaussNewtonRLforMPC.